# Natasha Kirova    Serguei Brazovskii

### v1

@@ Abstract
Phase transformations induced by short optical pulses is a mainstream in studies of of dynamics of cooperative electronic states. We present a semi-phenomenological modeling of spacio-temporal effects expected when optical excitons are intricate with the order parameter as it happens e.g. in organic compounds with neutral-ionic ferroelectric phase transitions. A conceptual complication appears here, where both the excitation and the ground sate ordering are built from the intermolecular electronic transfer. To describe both thermodynamic and dynamic effects on the same root, we adopt for the phase transition a view of the Excitonic Insulator - a hypothetical phase of a semiconductor which appears if the exciton energy becomes negative. After the initial pumping pulse, a quasi-condensate of excitons can appear as a macroscopic quantum state which then evolves interacting with other degrees of freedom prone to an instability. The self-trapping of excitons enhances their density which can locally surpass a critical value to trigger the phase transformation. The system is stratified in domains which evolve through dynamical phase transitions and may persist even after the initiating excitons have recombined.
@@ 1 Introduction
There is a persistent interest in phase transformations induced by optical pulses in cooperative electronic states (see e.g. a collection of articles [1], proceedings [2, 3] and the information from the latest conference [4]). The traditional technique of pump-induced phase transitions (PIPT) is based on the direct pumping between filled and empty electronic bands (the photon absorption energy $E_{ph}$ is above their separating gap $E_{g}$), creating the $e-h$ plasma to affect the correlated electronic states such as charge density waves (CDW), superconductors, or Peierls, Mott and excitonic insulators. The femtosecond shortness of pumps allows to resolve still disentangled ”fast” (10-100 fs) electronic and ”slow” (ps) lattice degrees of freedom. The intensity of pumps permits to kick the system far from the equilibrium which allows to follow the dynamical evolution over a wide range of the known phase diagram, see e.g. [5]. It becomes possible to reach also short-leaving ”hidden states” unavailable in thermodynamic approaches and even to obtain a truly stable one, see [6, 7].
@@
The exploitation of generically insulating systems, with their neutral excitations - the excitons, is a still ill explored land. The field is unlimited since many insulators are close to electronic and/or structural or magnetic instabilities, and all non-metallic systems possess some excitons accessible for pumping. The sufficiently dense gas of such excitons may be ready to form a coherent quantum state close to the Bose–Einstein condensation (BEC) [8]. Thus, the PIPT science for insulators needs also to include studies of coherent effects in ensembles of excitons. The BEC of excitons was searched for in bulk semiconductors, particularly in $Cu_{2}O$ [9] and finally observed in bilayer heterostructures. The BEC of excitons in the bulk is restricted to a low temperature $T_{BEC}$ because of their typically low concentration $n_{ex}$ limited by a ”short” $\backsim 10^{-9}$sec. life time. This is why the BEC was realized only for much lighter particles - the polaritons which entangle the exciton and the photon (see [10] for a review). A successful escape from this limitation has come with design of artificial bilayer heterostructures where electrons and holes are space separated, see [11] or e.g. [12]. We warn that PIPT offer another option: in conditions of the femtosecond pumping and the picoseconds’ scale of observation, $n_{ex}$ grows to huge levels of about one excitation for a few unit cells, yielding expectations for $T_{BEC}$ up to $100K$. The ”short” life time becomes ”long” more than enough.
@@
With a good luck, the pumping can go directly to the excitonic level by a resonant photon absorption $E_{ph}\approx E_{ex}$; then the excitations appear initially as delocalized plane waves, since photons create them at zero momentum $k=0$. In more common cases of a deviation from the resonance $E_{ph}>E_{ex}$, the hot excitons appear, still as delocalized waves, but with a distribution of momenta $k$, so that the phonon-assistant and/or collision processes are required for the energy relaxation. Cooling below the bandwidth of the excitonic spectrum $E_{ex}(k)$ towards its bottom (supposed to be at $k=0$) should lead to development of a quasi-condensate - the distribution peak at the lowest energy. The development of the BEC from the initially incoherent state is an intricate process, nowadays observed in cold atoms with many attempts to build a theory at both phenomenological [13] and microscopic [14, 15, 16] levels. The picture is not universal, depending on interactions of bosons (the pure repulsion for dipole excitons in double-wall structures, or the common, for excitons in the bulk, crossover among the attraction at large distances and the repulsion at short ones), and on inelastic scattering mechanisms required for cooling.
@@
Almost at the same time as Keldysh suggested the idea about the BEC of excitons, the concept of an ”excitonic insulator” (EI) as a special phase of a semiconductor was introduced [17, 18] after a vague proposal in [19] and its first theoretical development in [20]. The EI became a standard name for a state formed by the condensate of bound electron–hole pairs, on top of a semiconducting or a semimetallic state (see basic reviews [21, 22] and more modern ones in the collection [23]). The state of the EI appears when the total energy of the exciton vanishes, $E_{ex}=E_{g}-E_{b}\rightarrow 0$. That implies the possibility of independent manipulation of the conduction gap $E_{g}$ and the exciton’s binding energy $E_{b}$ (e.g., by the composition or the pressure). There is some confusion on the terminology since recently the name EI was assigned to states of a kind of CDWs [24, 25, 26, 27] where the lattice component is not seen, so the CDW is due to exclusively to $2P_{F}$ modulations of the electronic density; there can be no excitons with the well screened Coulomb interactions.
@@
Already in static conditions, the microscopic theory of the thermodynamic EI state and the theory of the BEC for optically pumped excitons are closely related, differing mostly by the monitoring parameters: the chemical potential $\mu_{ex}$ and the density $n_{ex}$ of excitons correspondingly. For the continuous optical pumping, this duality was strongly emphasized later [28].
@@
A dynamical theory of PIPTs is challenging for ab initio microscopic approaches, e.g. [29, 30, 31]. Nevertheless, the evolution at longer time scales is controlled by collective variables, such as the order parameter, excitons’ concentration and lattice displacements. Then a phenomenological approach grant us the possibility to describe oscillatory and inhomogeneous regimes of a dynamical symmetry breaking, already proved to be in a good accordance with experiment (see e.g. [5]). Another example was the modeling [32] of a stable hidden state of a polaronic Mott insulator observed in $1T-TaS_{2}$ [7]. The phenomenological approach becomes inevitable when we consider time dependent spacially inhomogeneous regimes which finally emerge in modeling.
@@
In this article we study the interplay among the pumping to excitons and the developing thermodynamic instabilities resulting in an induced phase transition. We present the phenomenological model which is based on a generalized concept of the EI formation evolving gradually from the BEC of excitons. On this bases we perform the modeling of spacio–temporal effects expected for the optically pumped ensemble of excitons coupled to the order parameter and lattice displacements. We shall briefly consider the case of extrinsic excitons which nature is essentially different with respect to the thermodynamical order parameter; the presence of excitons just shifts the transition temperature and the equilibrium value of the order parameters. Mainly we shall be concerned with a conceptually interesting case of intrinsic excitons which density contributes directly to the thermodynamic order parameter; this is the case of the charge transfer excitons with respect to the charge ordering transition. Keeping in mind the already existing experiments involving neutron-ionic transitions, we shall assume the second order character of the thermodynamic transition and include also possible interactions with another degree of freedom - the lattice displacements.
@@
We shall report two major observations. 1. The interaction of excitons with the order parameter can lead to their self-trapping enhances their local density which can locally surpass a critical value to trigger the phase transformation. The system is stratified in domains which evolve through dynamic phase transitions and may persist even after recombination of pumped excitons. 2. The excitons (unlike well studied cold Bose atoms) allow for virtual processes of creation/annihilation of their pairs from/to the vacuum which lead to a dynamical transition of the phase locking, thus indicating the demarcation between regimes of BEC and the EI. 3. We shall model some effects of transformations among the condensed and incoherent reservoirs of excitons. First, we shall study stochastic equations emulating a random falling of hot particles to the condensate. Second, we shall model a two-fluid macroscopic coexistence with a gradual absorption of the initially pumped particles to the condensate. Preliminary results and discussions can be found in earlier authors’ publications [33, 34, 35, 36].
@@
The article is organized as follows: Section 2 is alone devoted to case when the excitons are foreign with the order parameter. In Section 3 we the intricate case of the EI and the BEC. In Section 4 we present multi-field model for EI complemented by lattice displacements. Section 5 is dedicated to the effects coming from initially present hot excitons. The Section 6 is devoted to conclusions.
2 Extrinsic excitons and an arbitrary order parameter.
@@
Here and throughout the article, our basic assumption is that after a high initial pumping the quasi-condensate of optically pumped excitons appears sufficiently early as the macroscopic quantum state. The exciton is a quantum state in both the internal and the center of mass coordinates The condensate can be described by a macroscopic wave function $\Psi=|\Psi|\exp(i\zeta)$ which evolves interacting with other degrees of freedom prone to instability. The attempt to describe a coexistence of the BEC of excitons and their initially created incoherent liquid will be given below in Ch.5.
@@
$W(\eta)$ $=\frac{a}{2}\eta^{2}+\frac{g_{3}}{3}\eta^{3}+b\eta^{4}+\frac{c}{2}\left(\nabla\eta\right)^{2}$ (2)
Here the odd in $\eta$ terms (the ones with coefficients $g_{1,3}$) are present if the order parameter is not symmetry breaking which is commonly a case of the Ist order transition close to the IInd order one (the critical point of monitoring parameters where $a=0$). The odd $\varpropto g_{3}\eta^{3}$ term with $g_{3}<0$ can give rise to the the second, above the first one at $\eta_{1}=0$, minimum at some $\eta_{2}>0$ providing the appearance of the metastable state which further on becomes the preferable one. At presence of excitons with a density $n_{ex}=|\Psi|^{2}$, their interaction with $\eta$ shifts the critical point from $a_{cr}=0$ to $a_{cr}^{{}^{\prime}}=-g_{2}n_{ex}$. Reciprocally, the appearance of $\eta\neq 0$ shifts the exciton energy from $E_{ex}^{0}$ to $E_{ex}^{0}+g_{2}\eta^{2}/2$. The most common situation in PIPT, at least with conventional mechanisms of pumping to e-h excitations, is that the pumping destroys temporarily the low temperature (T) phase which corresponds to increase of $a$ from the negative value $a<0$ to the positive one $a^{\prime}=a+g_{2}n_{ex}>0$, hence expecting $g_{2}>0$. That can be verified by observing the increasing value of the exciton level in the ordered state $\eta\neq 0$. This situation is not a dogma and the opposite sign $g_{2}<0$ may take place. Intuitively, the sign of $g_{2}$ is expected to be positive if $\eta$ drives the system towards further dielectrization and negative if the system is driven towards metallization.
Within PIPT preferences, the most interesting scenario is a reaching of the second (meta)stable state at $\eta_{2}\neq 0$ which is imperative for a subset of non-symmetry-breaking transitions. That expectation does not exclude completely some rare symmetry breaking case where the negative higher order therms, like the change of sign in the coefficient $k$, are known to occur. In these cases, before pumping or at the low one, the phase with $\eta=0$ is stable and the phase with $\eta=\eta_{2}$ is unstable. As the result of pumping, the barrier and both energy minima diminish their values shifting at the same time their positions to higher $\eta$.
@@
Here $W$ is given by (1), $\omega$ is the bare frequency of oscillations of a weakly perturbed field $\eta$, $\gamma$ is their attenuation parameter, $\Gamma$ is the decay rate in the ensemble of excitons, $V$ is the system volume. Above, we have ignored effects breaking the conservation of the total number $N$ of excitons which will be important later for the case of intrinsic excitons, in Section 3.
The first critical pumping $n_{ex}=n_{c}$ corresponds to the binodal formation: the energies of both energy minima are equal, being still separated by the barrier. The second critical pumping at $n_{ex}=n_{d}$ corresponds to the instability of the phase 1 when the barrier disappears. At $n_{ex}\geq n_{d}$ the phase 1 is metastable; at the even higher pumping $n_{ex}\geq n_{d}$ the energy minimum $W_{1}(\eta)$ drops below the initial $W(0)=0$.
After the instantaneous switching, the excess potential energy transforms into the kinetic energy giving rise to collective pendulum oscillations. After a number of oscillations, the system achieves the minimum corresponding to the pure dynamical transition when $W_{b}(\eta_{d})=0$. The state 1 becomes metastable as the state 2 is well below in energy. Fluctuations of basic parameters result also in formation of alternating domains of coexisting phases. The third critical pumping appears when the spinodal is formed, the barrier merges with the phase 1 minimum which corresponds to the absolute instability of the phase 1. The pumping threshold $n_{c3}$ does not dependent on excitons’ coupling parameters.
The numerical modeling of a similar system can be found in [33].
3 Intrinsic excitons and the excitonic insulator.
3.1 Generic excitonic insulator.

### v2

# Inhomogeneous Evolution of a Dense Ensemble of Optically Pumped Excitons to a Charge Transfer State

@@ Abstract
Phase transformations induced by short optical pulses are mainstream in studies on the dynamics of cooperative electronic states. We present a semi-phenomenological modeling of spacio-temporal effects expected when optical excitons are intricate with the order parameter as in, e.g., organic compounds with neutral-ionic ferroelectric phase transitions. A conceptual complication appears here, where both the excitation and the ground state ordering are built from the intermolecular electronic transfer. To describe both thermodynamic and dynamic effects on the same root, we adopt, for the phase transition, a view of the Excitonic Insulator—a hypothetical phase of a semiconductor that appears if the exciton energy becomes negative. After the initial pumping pulse, a quasi-condensate of excitons can appear as a macroscopic quantum state that then evolves, while interacting with other degrees of freedom which are prone to an instability. The self-trapping of excitons enhances their density, which can locally surpass a critical value to trigger the phase transformation. The system is stratified in domains that evolve through dynamical phase transitions and may persist even after the initiating excitons have recombined.
Keywords: excitonic insulator; exciton; optical pumping; PIPT; BEC; self-trapping
@@ 1 Introduction
There is a persistent interest in phase transformations induced by optical pulses in cooperative electronic states (see, e.g., a collection of articles [1], proceedings [2, 3], and information from the latest conference [4]). The traditional technique of pump-induced phase transitions (PIPT) is based on the direct excitation between filled and empty electronic bands (the photon absorption energy $E_{ph}$ is above their separating gap $E_{g}$), creating $e-h$ plasma to affect the correlated electronic states, such as charge density waves (CDWs), superconductors, or Peierls, Mott, and excitonic insulators. The shortness of pumps (of the order of femtoseconds) allows one to resolve still disentangled “fast” (10–100 fs) electronic and “slow” (ps) lattice degrees of freedom. The intensity of the pumps allows for the system to be kicked far away from the equilibrium, which allows one to follow the dynamical evolution over a wide range of the known phase diagram; see e.g., [5]. It becomes possible to also reach short-leaving “hidden states” unavailable in thermodynamic approaches and even to obtain a truly stable state; see [6, 7].
@@
The exploitation of generic insulating systems, with their neutral excitations—the excitons—is still an ill-explored land. The field is unlimited since many insulators are close to electronic, structural, and/or magnetic instabilities, and all non-metallic systems possess some excitons accessible for pumping. The sufficiently dense gas of such excitons may be ready to form a coherent quantum state close to the Bose–Einstein condensation (BEC) [8]. Thus, the PIPT science for insulators needs to also include studies on the coherent effects in ensembles of excitons. The BEC of excitons was searched for in bulk semiconductors, particularly in $Cu_{2}O$ [9], and finally observed in bi-layer heterostructures. The BEC of excitons in the bulk is restricted to a low-temperature $T_{BEC}$ because of their typically low concentration $n_{ex}$ limited by a ”short” $\backsim 10^{-9}$ s life time. This is why the BEC was realized only for much lighter particles—the polaritons that entangle the exciton and the photon (see [10] for a review). This limitation was successfully avoided with a design of artificial bi-layer heterostructures, where electrons and holes are space-separated; see short reviews in [11] or, e.g., [12].
We propose that PIPTs offer another option: under the conditions of femtosecond pumping and the picoseconds scale of observation, $n_{ex}$ increases to high levels of about one excitation for a few unit cells, yielding expectations for $T_{BEC}$ up to $100$ K. The “short” life time becomes more than “long” enough.
@@
With good luck, the pumping can go directly to the excitonic level using a resonant photon absorption $E_{ph}\approx E_{ex}$; then, the excitations appear initially as delocalized plane waves, since photons create them at zero momentum $k=0$. In more common cases of a deviation from the resonance $E_{ph}>E_{ex}$, hot excitons appear, still as delocalized waves but with a distribution of momenta $k$, so that phonon assistant and/or collision proOKcesses are required for energy relaxation. Cooling below the bandwidth of the excitonic spectrum $E_{ex}(k)$ toward its bottom (which is supposed to be at $k=0$) should lead to the development of a quasi-condensate—the distribution peak at the lowest energy. The development of the BEC from the initially incoherent state is an intricate process, having encouraged many attempts at building a theory at both phenomenological [13] and microscopic [14, 15, 16] levels. The resulting picture is not universal as it depends on the interactions of bosons (either pure repulsion for dipole excitons in double-wall structures, or the common for excitons crossover from attraction at large distances to repulsion at short ones) and on the inelastic scattering mechanisms required for cooling.
@@
Almost at the same time that Keldysh suggested the idea of the BEC of excitons, the concept of an “excitonic insulator” (EI) as a special phase of a semiconductor was introduced [17, 18] after a vague proposal in [19] and its first theoretical development in [20]. The EI became a standard name for a state formed by the condensate of bound electron–hole pairs, on top of a semiconducting or a semi-metallic state (see basic reviews [21, 22] and more modern ones in collection [23]). The state of the EI appears when the total energy of the exciton vanishes, $E_{ex}=E_{g}-E_{b}\rightarrow 0$. This implies the possibility of the independent manipulation of the conduction gap $E_{g}$ and the exciton binding energy $E_{b}$ (e.g., by the composition or the pressure or the bias voltage). There is some confusion on the terminology since, recently, the name EI was assigned to states of a kind of CDWs [24, 25, 26, 27] where the lattice component is not seen, so the CDW is exclusively caused by $2P_{F}$ modulations of the electronic density; there can be no excitons with well-screened Coulomb interactions.
@@
In static conditions, the microscopic theory of the thermodynamic EI state and the theory of the BEC for optically pumped excitons are closely related, differing mostly in their monitoring parameters: the chemical potential $\mu_{ex}$ and the density $n_{ex}$ of excitons, respectively. For continuous optical pumping, this duality was strongly emphasized in [28].
@@
A dynamical theory of PIPTs is challenging for ab initio microscopic approaches; see, e.g., [29, 30, 31]. Nevertheless, evolution at longer time scales is controlled with collective variables, such as order parameter, the excitons’ concentration, and lattice displacements. Then, a phenomenological approach grants us the possibility to describe oscillatory and inhomogeneous regimes of dynamical symmetry breaking, which has already been proved to be in good accordance with experimentation (see, e.g., [5]). Another example was the modeling [32] of a stable hidden state of a polaronic Mott insulator observed in $1T-TaS_{2}$ [7]. The phenomenological approach becomes inevitable when we consider time dependent spacially inhomogeneous regimes which finally emerge in modeling.
@@
In this article we study the interplay among the pumping to excitons and the developing thermodynamic instabilities resulting in an induced phase transition. We present the phenomenological model which is based on a generalized concept of the EI formation evolving gradually from the BEC of excitons. On this bases we perform the modeling of spacio–temporal effects expected for the optically pumped ensemble of excitons coupled to the order parameter and lattice displacements. We shall briefly consider the case of extrinsic excitons which nature is essentially different with respect to the thermodynamical order parameter; the presence of such excitons just shifts the transition temperature and the equilibrium value of the order parameters. Mainly we shall be concerned with a conceptually most interesting case of intrinsic excitons which density contributes directly to the thermodynamic order parameter; this is the case of the charge transfer excitons with respect to the charge ordering transition. Keeping in mind the already existing experiments involving neutron-ionic ferroelectric transitions, we shall include also possible interactions with another degree of freedom—the lattice displacements.
@@
We shall report on two major observations. 1. The interaction of excitons with the order parameter can lead to their self-trapping thus enhancing their local density which can locally surpass a critical value to trigger the phase transformation. Then the system is stratified in domains which evolve through dynamical phase transitions and may persist even after recombination of pumped excitons. 2. The excitons (unlike well studied cold Bose atoms) allow for virtual processes of creation/annihilation of their pairs from/to the vacuum which lead to a dynamical transition of the phase locking, thus indicating the demarcation between regimes of BEC and the EI.
We model some effects of transformations among the condensed and incoherent reservoirs of excitons. First, we shall study stochastic equations emulating a random falling of hot particles to the condensate. Second, we shall model a two-fluid macroscopic coexistence with a gradual absorption of the initially pumped particles to the condensate. Preliminary results and discussions can be found in earlier authors’ publications [33, 34, 35].
@@
The article is organized as follows: Section 2 is alone devoted to the case when the excitons are foreign with respect to the order parameter. In Section 3 we introduce the intricate case of the EI and the BEC. In Section 4 we present multi-field model for the EI complemented by lattice displacements. Section 5 is dedicated to the effects coming from initially present hot excitons. Section 6 is devoted to conclusions.
2 Extrinsic Excitons and an Arbitrary Order Parameter
@@
The single exciton is a quantum state in both the internal and the center of mass coordinates. Here and throughout the article, our basic assumption will be that, after a high initial pumping, the quasi-condensate of optically pumped excitons appears sufficiently early as the macroscopic quantum state. The condensate can be described by a macroscopic wave function $\Psi=|\Psi|\exp(i\zeta)$ which evolves interacting with other degrees of freedom prone to an instability. The attempt to describe a coexistence of the BEC of excitons and their initially created incoherent liquid will be given below in Section 5.
@@
$W(\eta)=\frac{a}{2}\eta^{2}+\frac{g_{3}}{3}\eta^{3}+\frac{b}{4}\eta^{4}+\frac{c}{2}\left(\nabla\eta\right)^{2}$
Here the odd in $\eta$ terms (the ones with coefficients $g_{1,3}$) are present if the order parameter is not symmetry breaking which is commonly a case of the Ist order transition close to the IInd order one (the critical point of monitoring parameters where $a=0$). The odd $\varpropto g_{3}\eta^{3}$ term with $g_{3}<0$ can give rise to the second, above the first one at $\eta_{1}=0$, minimum at some $\eta_{2}>0$ bringing to life the metastable state which further on becomes the preferable one. At presence of excitons with a density $n_{ex}=|\Psi|^{2}$, their interaction with $\eta$ shifts the critical point from $a_{cr}=0$ to $a_{cr}^{{}^{\prime}}=-g_{2}n_{ex}$. Reciprocally, the appearance of $\eta\neq 0$ shifts the exciton energy from $E_{ex}^{0}$ to $E_{ex}^{0}+g_{2}\eta^{2}/2$. That can be verified by observing the increasing value of the exciton level in the ordered state $\eta\neq 0$. This situation is not a dogma and the opposite sign $g_{2}<0$ may take place. Intuitively, the sign of $g_{2}$ is expected to be positive if $\eta$ drives the system towards further dielectrization and negative if the system is driven towards metallization.
@@
Here $W$ is given by Equation (1), $\omega$ is the bare frequency of oscillations of a weakly perturbed field $\eta$, $\gamma$ is their attenuation parameter, $\Gamma$ is the decay rate in the ensemble of excitons, $V$ is the system volume. Above, we have ignored effects breaking the conservation of the total number $N$ of excitons which will be important later for the case of intrinsic excitons, in Section 3.
[…]
With the later PIPT tendency of searching for “hidden states”, the preferable scenario became to reach the second (meta)stable state at some $\eta_{2}\neq 0$. Its existence is imperative for non-symmetry-breaking transitions where $g_{3}\neq 0$. That does not exclude some rare symmetry breaking cases where the negative higher order therms, like the coefficient $b$, are known to occur, leading also to a weakly Ist order phase transition. In these cases, at high T the lowest state 1 with $\eta=0$ is stable and the higher energy state 2 with $\eta=\eta_{2}$ is metastable; the states are separated by a barrier at some $\eta=\eta_{b}$. With decreasing $a$ the binodal is reached when both states 1 and 2 are of the same energy. If the state 1 survives as now the metastable state, than a spinodal point $a=0$, with no more barrier, can be reached in similarity to the above case of the IInd order transition.
Now two directions of PIPT are envisionable with different types of switching: i. from the high T state 1 to the state 2 provided that $g_{2}<0$ or ii. from the low T state 2 to the state 1 provided that $g_{2}>0$. The numerical modeling of a kind of such a system can be found in [33]. An intrigue was that in a certain interval around the barrier $\eta_{b}$, the energy as a function of $\eta$ has a negative curvature bringing to life the space homogeneous regime with rich patterns of arising, colliding, and merging domains and their walls. Beyond the conventional in thermodynamics picture of the separation of coexisting phases, in PIPT dynamics these patterns serve also as self-trapping areas for excitons where their local density is enhanced with respect to the mean one, then the induced transitions take place locally even before the mean value of the critical pumping is reached for the whole volume.
3 Intrinsic Excitons and the Excitonic Insulator
3.1 Generic Excitonic Insulator